OpenStudy (anonymous):

Could someone help me with this please? Variable y varies directly with x2, and y = 72 when x = 6. Enter the constant of variation for the quadratic variation equation. y = ? x2 I think the answer is 2 but I'm not very sure

OpenStudy (cwrw238):

y = kx^2 where k is the constant of variation if you plug in x = 6 and y = 72 into the above you can calculate k;- 72 = k * (6)^2

OpenStudy (cwrw238):

so 36k = 72 now divide both sides by 36

OpenStudy (cwrw238):

so what is the value of k?

OpenStudy (anonymous):

72/36 would be 2 so would k = 2?

OpenStudy (holly00d1248):

I agree with you @--Irwin